OpenStudy (anonymous):

The greatest common factor of 15 and 45 is?

OpenStudy (anonymous):

try using 5. because 5(3) is 15 and 5(9) is45

OpenStudy (anonymous):

look at prime factorisation: 15=3*5 45=3^2*5 so mutual primes are 3 and 5 and GCF is 5*3 = 15 I'm 100% sure and try the Maths section next time :)